# Depth-Based Local Center Clustering: A Framework for Handling Different Clustering Scenarios

### v1

@@ V Experiments and Evaluations
External Metric. To compare the performance of different clustering algorithms on the same datasets, we use the clustering error (CE) [28, p. 2774] and adjusted Rand index (ARI) [51]. The CE simply computes the proportion of mismatched points. This makes CE intuitive, with a value range of 00 (perfect match) to $1$. The ARI, on the other hand, measures pairwise similarity between clustering results and ground truth. The expected value fo the ARI under random class assignment is 0 and it takes a value $1$ under perfect agreement.
Methods for comparison. We compare our approach with several representative clustering methods. These include the center-based algorithm K-means and its state-of-the-art variant LW-K-means; the recent advanced DPC method R-MDPC; the subspace clustering algorithm SSC; the model-based clustering algorithm GMM (GPCM family);
V-A Description of Datasets
Figure 7 presents four synthetic datasets. The first subfigure is simulated using custom code, while the second and third subfigures are generated with the mlbench [53] package in R [52]. The last subfigure provides a two-dimensional visualization of the ray shapes data, which is derived from a study on spectral clustering [54]. These synthetic datasets represent different clustering scenarios with non-convex shapes, clusters of varying sizes and densities in close proximity, and large overlapping regions. Each of these characteristics presents unique challenges for clustering.
Figure 7: Graphs of synthetic datasets: (a) Starbeam; (b) Blend; (c) Bainba; (d) Ray [54]. In $(d)$, $x$ and $y$ axes are the first two variables of ray data. Colors represent true labels.
As for the real-world datasets, we utilize five benchmark datasets from the UCI Machine Learning Repository [55]: iris, seed, wine, optidigits, pendigits. Additionally, we test high-dimensional datasets ($d\geq 100$) with $n>d$, including Yale B [56] and HAR [57], which represent human-face and human-activity datasets, respectively, addressing clustering problems in different domains. Among these, Yale B, optidigits, and pendigit are sourced from the R package PPCI [58]. Furthermore, we consider two microarray datasets: leukemia [59] and SuCancer [60], which have large dimensions $d$ but small sample sizes $n$.
V-B Result comparisons
Table IV provides the settings, used for running DLCC on the above-mentioned datasets. The max strategy tends to perform better with non-convex or unbalanced sized clusters. Conversely, the min strategy is more effective for overlapping clusters of similar sizes.
TABLE IV: DLCC settings for different datasets, where CM stands for classification method.
Datasets Depth Strategy CM maxdepth ifloop
Starbeam MD Max KNN False --
Starbeam SD Max KNN False --
Blend MD Max KNN False --
Blend SD Max KNN False --
Bainba MD Max MDC True --
Bainba SD Max KNN False --
Ray MD Min MDC False True
Ray SD Min RF False False
Iris MD Min MDC True False
Iris SD Min RF False False
Seed MD Min KNN True True
Seed SD Min KNN False False
Wine MD Min MDC True True
Wine SD Min MDC False False
Yale B SD Min RF False False
Optidigit SD Min KNN False False
Pendigits MD Min KNN False False
Pendigits SD Min KNN False False
HAR SD Min KNN False False
Leukemia SD Max MDC False --
SuCancer SD Max RF False --
@@ V-B Result comparisons
TABLE V: Comparative analysis of selected clustering algorithms on several datasets ( “–” means a method is not applicable).
Datasets Metric DLCCMD DLCCSD Kmeans LW-Kmeans SSC GMM R-MDPC
Starbeam ARI $\mathbf{1}$ $\mathbf{1}$ $0.4822$ $0.4822$ $0.3052$ $0.0157$ $0.8123$
$n=780,d=2,K=2$ CE $\mathbf{0}$ $\mathbf{0}$ $15.26\%$ $15.26\%$ $22.31\%$ $40.26\%$ $4.87$%
Blend ARI 0.85950.8595{0.8595}0.8595 $\mathbf{0.8605}$ $0.7266$ $0.7365$ $0.5716$ $0.7518$ $0.8453$
$n=2000,d=2,K=10$ CE ${12.50}\%$ $\mathbf{8.65}\%$ $25.70\%$ $18.90\%$ $31.35\%$ $21.45\%$ ${9.10}\%$
Bainba ARI $\mathbf{1}$ $\mathbf{1}$ $0.4961$ $0.9814$ $0.4276$ $0.7436$ $\mathbf{1}$
$n=600,d=2,K=3$ CE $\mathbf{0}$ $\mathbf{0}$ $32.00\%$ $0.67\%$ $30.17\%$ $10.50\%$ $\mathbf{0}$
Ray ARI $\mathbf{0.8261}$ $0.7521$ $0.6388$ $0.7090$ $0.7081$ $0.8245$ $0.5364$
$n=1000,d=5,K=4$ CE $\mathbf{7.00}\%$ $10.00\%$ $15.00\%$ $11.90\%$ $12.60\%$ $7.10\%$ $37.70\%$
Iris ARI $\mathbf{0.9039}$ $0.8857$ $0.7302$ $0.8857$ $0.6138$ $\mathbf{0.9039}$ $0.8857$
$n=150,d=4,K=3$ CE $\mathbf{3.33}\%$ $4.00\%$ $10.67\%$ $4.00\%$ $17.33\%$ $\mathbf{3.33}\%$ $4.00\%$
Seed ARI $0.7612$ $\mathbf{0.7626}$ $0.7167$ $0.6687$ $0.4462$ $0.6299$ $0.7289$
$n=210,d=7,K=3$ CE $\mathbf{8.57}\%$ $\mathbf{8.57}\%$ $10.48\%$ $12.86\%$ $33.33\%$ $14.76\%$ $10.00\%$
Wine ARI $\mathbf{0.9817}$ $0.9295$ $0.8975$ $0.8516$ $0.8804$ $0.9309$ $0.7269$
$n=178,d=13,K=3$ CE $\mathbf{0.56}\%$ $2.25\%$ $3.37\%$ $5.06\%$ $3.93\%$ $2.25\%$ $9.66\%$
Yale B ARI -- $\mathbf{0.9901}$ $0.5663$ $0.8461$ $0.9484$- $0.7635$
$n=2000,d=600,K=10$ CE -- $\mathbf{0.45}\%$ $31.85\%$ $13.25\%$ $2.40\%$- $12.10\%$
Optidigits ARI -- $0.8144$ $0.5682$ $0.5799$ $0.7540$ $0.4878$ $\mathbf{0.8543}$
$n=5620,d=64,K=10$ CE -- $9.40\%$ $28.81\%$ $28.26\%$ $14.84\%$ $36.89\%$ $\mathbf{8.75}\%$
Pendigits ARI $0.6488$ $0.6661$ $0.5526$ $0.5756$ $\mathbf{0.7069}$ $0.6522$ $0.6657$
$n=10992,d=16,K=10$ CE $20.52\%$ $18.22\%$ $32.25\%$ $27.83\%$ $\mathbf{15.56}\%$ $20.58\%$ $19.17\%$
HAR ARI -- $0.6564$ $0.4610$ $0.4556$ $\mathbf{0.7005}$- $0.4373$
$n=10299,d=561,K=6$ CE -- $\mathbf{18.68}\%$ $40.00\%$ $41.51\%$ $23.11\%$- $51.06\%$
Leukemia ARI -- $\mathbf{0.8897}$ $\mathbf{0.8897}$ $\mathbf{0.8897}$ $0.7371$- $0.7851$
$n=72,d=3571,K=2$ CE -- $\mathbf{2.78}\%$ $\mathbf{2.78}\%$ $\mathbf{2.78}\%$ $6.94\%$- $5.56\%$
SuCancer ARI -- $0.1154$ $-0.003$ $-0.003$ $\mathbf{0.1234}$- $0.1154$
$n=174,d=7909,K=2$ CE -- $32.76\%$ $47.70\%$ $47.70\%$ $\mathbf{32.18}\%$- $32.76\%$
Table V illustrates the performance of DLCC with both MD and SD, compared to other clustering algorithms. For all comparing methods, we pick the result achieving the highest ARI as the best. For K-means and LW-K-means, we report the best results from $10$ runs. For the R-MDPC method, following the suggestions, we fix $cv=1$ and test all neighborhood sizes from $0.5\sqrt{n}$ to $2\sqrt{n}$ [18]. For SSC, the best combinations are chosen with regards to the hyper parameter $\alpha$ (ranging from $10$ to $50$ with a step size of $10$) and the logical parameters “affine” (indicating if affine subspaces are considered) and “outlier” (signifying if the data might be contaminated by outliers). In the case of GMM, the R package mixture [61] is used to identify the optimal model from the GPCM family. As for DLCC, when RF is employed as the classification method, its inherent randomness led us to conduct $100$ separate runs with unique seeds. The corresponding internal metric for each run was calculated, and the median performance score was recorded in the table.
Table V clearly demonstrates that DLCC yields impressively competitive results across all datasets. Specifically, in the synthetic datasets, both versions of DLCC perform perfectly in Starbeam and Bainba, while they meet some challenges in Blend. As detailed in Section III-A, the max strategy assumes that, within any given cluster, each subset used for defining local centers is convex. Although the filtering procedures for local centers can identify the areas of non-convex subsets, they are unable to further distinguish individual clusters within these regions and, in the current context, to identify the two distinct spirals. DLCCMD achieves the best performance on the Ray dataset, closely matching that of GMM, as its similarity matrix is constructed based on the EEV model, a member of GMM. DLCCSD proves highly adaptable for high-dimensional data applications, delivering the best results in the Yale B and Leukemia datasets and ranking among the top performers on other high-dimensional datasets. In contrast, each of the other algorithms displays distinct strengths and limitations. For instance, K-means and LW-Kmeans produce satisfactory results for convex datasets, while they do not perform well on datasets with non-convex shapes. R-MDPC performs well on non-convex datasets and excels on the Optidigits dataset. However, it faces challenges with datasets that have large overlapping regions with similar densities, such as Ray and HAR. SSC excels with high-dimensional datasets but falters when the dimension $d$ is less than or close to the number of clusters $K$. Both GMM and DLCCMD face challenges in high-dimensional data due to their reliance on covariance matrix estimations. Accurately estimating covariance matrices requires significantly more observations per cluster than the number of dimensions, which is often not feasible for high-dimensional data.
V-C Performance in estimating the number of cluster
For the max strategy, the relationship between $\delta$ and the number of clusters is detailed in Section IV. Here, we evaluate the effectiveness of Algorithm 1 in estimating the number of clusters under the min strategy. To ensure a fair comparison, the results presented in Table V assume that the true number of clusters is known. In this section, we focus on the scenario where $K$ is not pre-specified, specifically discussing DLCCSD as it is applicable to all datasets.
To provide insights into the performance of our method, we compare it with a well-established spectral clustering method, Spectrum [62], which can automatically estimate the number of clusters during the clustering process without requiring extensive parameter tuning. The implementation of Spectrum is based on the R package Spectrum [63], and we report the better results between the eigengap method (largest gap between eigenvalues) and the multimodality gap method.
Table VI demonstrates that without pre-defined $K$, DLCCSD consistently outperforms Spectrum in terms of ARI. Although DLCCSD estimates two additional clusters in the Pendigits dataset, its ARI is even higher compared to the scenario where $K$ is fixed at $10$. For the Optidigits and HAR datasets, DLCCSD underestimates the number of clusters by one. Specifically, it merges clusters “3” and “8” into a single cluster in Optidigits, and in HAR, it groups walking, walking upstairs, and walking downstairs into one cluster. This clustering also makes sense based on human assessments.
TABLE VI: Comparison of performance between DLCCSD and Spectrum across datasets. Here, $\hat{K}$ denotes the estimated number of clusters, with differences from true $K$ in parentheses.
Method Ray Iris Seed Wine Yale B Optidigits Pendigits HAR
$\hat{K}$ ARI $\hat{K}$ ARI $\hat{K}$ ARI $\hat{K}$ ARI $\hat{K}$ ARI $\hat{K}$ ARI $\hat{K}$ ARI $\hat{K}$ ARI
DLCCSD $4$ (00) $0.7521$ $3$ (00) $0.8857$ $3$ (00) $0.7626$ $3$ (00) $0.9295$ $10$ (00) $0.9901$ $9$ ($-1$) $0.7174$ $12$ ($+2$) $0.6885$ $4$ ($-2$) $0.4813$
Spectrum $3$ ($-1$) $0.6468$ $3$ (00) $0.8508$ $3$ (00) $0.5702$ $3$ (00) $0.9170$ $9$ ($-1$) $0.7575$ $9$ ($-1$) $0.6527$ $10$ (00) $0.5086$ $10$ ($+4$) $0.3242$
V-D Complexity Analysis
The computational cost of the DLCC algorithm is dominated by constructing the depth-based similarity matrix and estimating local centers. For both versions, the first term corresponds to constructing the similarity matrix, and the second term accounts for estimating local centers, with $s\ll n$.
  • DLCCMD: With pre-estimated covariance matrices (Section II-B), the complexity is $O(n^{2})+O(ns^{2}+ns)$, where $O(ns^{2}+ns)$ covers covariance matrix estimation and depth computation for neighborhoods. As $s\ll n$, the overall complexity simplifies to $O(n^{2})$.
  • DLCCSD: The complexity is $O(n^{3})+O(ns^{2})$. Constructing the similarity matrix requires $O(n^{3})$, as computing SD values for $n$ points ($O(n^{2})$) across $n$ reflected datasets introduces an additional $n$ factor. Estimating local centers for neighborhoods adds $O(ns^{2})$, resulting in an overall complexity of $O(n^{3})$.
Figure 8: Running time of DLCC on simulated datasets without acceleration techniques, where blue represents DLCCSD and red represents DLCCMD: (a) Fixed the dimension at $30$ while increasing the sample size up to $5060$, y-axis up to $700$s; (b) Fixed the sample size at $1000$ while increasing the dimension up to $1000$, y-axis up to $25$s.
Figure 8(a) shows that the experimental running times align with theoretical complexities: $O(n^{2})$ for DLCCMD and $O(n^{3})$ for DLCCSD. In Figure 8(b), the running time of DLCCMD appears irregular as dimensionality increases because when estimating the covariance matrix, the method defaults to the identity matrix if singularity issues arise, which are more likely when the dimensionality is not significantly smaller than the sample size. In contrast, DLCCSD shows a linear trend due to the SD computation involving the $L^{2}$ norm of the average of unit vectors, with a complexity of $O(n^{3}d)$ when $d$ is large.
While DLCC can handle datasets with approximately $10,000$ samples, its computational cost is dominated by the depth-based similarity matrix, limiting its scalability to larger datasets. However, DLCCSD is computationally efficient with increasing dimensionality. Future work could explore approximation methods to accelerate the algorithm. For instance, defining local centers does not necessarily require computing the entire similarity matrix but rather identifying suitable neighborhoods for the centers. Fast approximate neighborhood search methods (e.g., [64]) could be employed to estimate candidate local centers (with size much smaller than $n$), followed by exact depth similarity computations for points relative to these candidates.
VI Conclusions and Future work

### v2

@@ V-A External Metric
To compare the performance of different clustering algorithms on the same datasets, we use the clustering error (CE; [57, p. 2774]), ARI and normalized mutual information (NMI; [58]). The CE simply computes the proportion of mismatched points. This makes CE intuitive, with a value range of 0 (perfect match) to $1$. The ARI, on the other hand, measures pairwise similarity between clustering results and ground truth. The expected value fo the ARI under random class assignment is 0 and it takes a value $1$ under perfect agreement. NMI measures the shared information between the predicted and true partitions and is normalized to $[0,1]$, with $1$ indicating perfect agreement and larger values implying better alignment.
V-B Methods for Comparison
We compare DLCC with representative clustering methods that rely on exemplars or explicitly group local sub-regions/exemplars which show resemblance of DLCC, including LW-k-means [12], EKM [14], R-MDPC [18], granular-ball clustering (GBCT; [20]), and ACLR [27]. We also include a recent depth-based alternative to mean-shift [35], denoted as DMS. For all those methods except DMS, we use the authors’ released implementations; DMS is implemented by us using lens depth [59] since no existing implementation can be found.
All benchmark datasets considered in this paper are equipped with ground-truth labels. Since a single solution is generally not expected to simultaneously optimize all three external metrics (ARI, NMI, and CE), we adopt ARI as the primary selection metric throughout. Whenever a method involves hyperparameter tuning, we select and report the solution with the best ARI, and report the corresponding NMI and CE for the same solution.
For fair comparisons, we assume the true number of clusters $K$ is known, and set $K$ accordingly for methods that require it. For LW-k-means, we report the median performance over $100$ random initializations. For EKM, we set the number of replications to $1000$ and report the solution with the smallest objective value (as in the original implementation). For R-MDPC, we follow the authors’ recommendation by fixing $cv=1$ and scanning neighborhood sizes in $0.5\sqrt{n},\,2\sqrt{n}]$. For ACLR, we scan the hierarchy of binary trees $h\in\{7,\ldots,11\}$ and the anchor-graph neighborhood size in $[10,100]$. For DMS, we scan the quantile level used to determine the localization parameter $\tau$ over $[0.01,0.5]$. The direction parameter is tuned as follows: we first evaluate two initial values ($20$ and $500$) and choose the better one as the starting point. We then perform a one-sided search by increasing from $20$ (or decreasing from $500$) and stop when the ARI drops by more than $0.05$ for two consecutive steps.
V-C Synthetic Datasets
We evaluate DLCC on $12$ synthetic datasets designed to cover a broad range of clustering scenarios and to give insights on both the strengths and limitations of the proposed DLCC algorithm. Figure 7 visualizes the DLCC partitions. The datasets 3 blobs, Aggregation, Atom, Chainlink, Cuboids, Diamond, Flame, and Jain are benchmark examples from [60]. Bainba and Blend are generated using the mlbench package [61] in R [62]. Ray is adapted from a spectral-clustering study [63], and Starbeam is constructed by us.
Following the order in Figure 7, the synthetic datasets represent the following clustering challenges:
  1. (a)
    well-separated, convex shapes.
  2. (b)
    bridge-connected shapes with varying sizes.
  3. (c)
    a non-convex component enclosing a convex core.
  4. (d)
    proximity-driven ambiguity under strong size imbalance.
  5. (e)
    arbitrary shapes with substantial heterogeneity.
  6. (f)
    non-convex interlocked rings.
  7. (g)
    well-separated shapes with different orientations.
  8. (h)
    close convex shapes.
  9. (i)
    close non-convex shapes with similar density.
  10. (j)
    two curved non-convex shapes separated by a narrow gap, with varying densities.
  11. (k)
    overlapping/entangled shapes.
  12. (l)
    non-convex connected shapes with varying densities.
In these experiments, we adopt the min strategy for 3 blobs, Diamond, and Ray, and the max strategy for the remaining datasets (refer to Section IV-A for reasons).
Figure 7: DLCC clustering visualizations on 12 synthetic datasets. For the Ray dataset ($d=5$), the plot is produced by projecting the data onto the first two principal components.
TABLE II: Comparative analysis of selected clustering algorithms on synthetic datasets (For methods without a fixed $K$, when the estimated $\hat{K}$ differs from the true $K$, the CE is not meaningful and we report $\hat{K}$ instead).
Datasets Metric DLCC LW-k-means EKM R-MDPC GBCT ACLR DMS
3 blobs $n=300,d=2,K=3$ ARI NMI CE $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $0.9900$ $0.9830$ $0.33\%$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $0.9900$ $0.9830$ $0.33\%$ $0.9900$ $0.9830$ $0.33\%$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$
Aggregation $n=788,d=2,K=7$ ARI NMI CE $0.9927$ $0.9883$ $0.38\%$ $0.7434$ $0.8409$ $13.07\%$ $0.6325$ $0.7646$ $29.31\%$ $0.8415$ $0.9091$ $14.09\%$ $0.7359$ $0.8378$ $21.57\%$ $\boldsymbol{0.9956}$ $\boldsymbol{0.9923}$ $\boldsymbol{0.25\%}$ $0.9109$ $0.9503$ $\hat{K}=6$
Atom $n=800,d=3,K=2$ ARI NMI CE $0.0193$ $0.1220$ $43.00\%$ $0.1929$ $0.3044$ $28.00\%$ $0.1066$ $0.2332$ $33.63\%$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $0.7587$ $0.8201$ $20.25\%$
Bainba $n=600,d=2,K=3$ ARI NMI CE $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $0.9814$ $0.9686$ $0.67\%$ $0.4600$ $0.5016$ $38.50\%$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $0.7578$ $0.7764$ $9.67\% 𝟏\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $0.9600$ $0.9362$ $\hat{K}=4$
Blend $n=2000,d=2,K=10$ ARI NMI CE $0.8579$ $0.9588$ $\hat{K}=9$ $0.7310$ $0.8432$ $19.58\%$ $0.5793$ $0.7363$ $32.50\%$ $0.8337$ $0.9448$ $16.35\%$ $0.9656$ $0.9795$ $45\% 𝟏\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $0.8218$ $0.8985$ $\hat{K}=14$
Chainlink $n=1000,d=3,K=2$ ARI NMI CE $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ 0.0010-0.0010 $0.0000$ $49.80\%$ 0.0010-0.0010 $0.0000$ $49.70\%$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$
Cuboids $n=1002,d=3,K=4$ ARI NMI CE $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}a$ $0.6713$ $0.7390$ $28.14\%$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$
Diamond $n=3000,d=2,K=9$ ARI NMI CE $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $0.7855$ $0.8600$ $16.33\%$ $0.9992$ $0.9990$ $0.03\%$ $0.8758$ $0.9473$ $0.27\% 𝟏\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$
Flame $n=240,d=2,K=2$ ARI NMI CE $0.9666$ $0.9269$ $0.83\%$ $0.4880$ $0.4422$ $15.00\%$ $0.3988$ $0.3544$ $8.33\% 𝟏\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $0.9501$ $0.8991$ $1.25\%$ $0.9666$ $0.9269$ $0.83\%$ $0.9666$ $0.9269$ $0.83\%$
Jain $n=373,d=2,K=2$ ARI NMI CE $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $0.6352$ $0.5908$ $9.92\%$ $0.5681$ $0.5088$ $2.06\% 𝟏\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $0.2563$ $0.2463$ $9.03\% 𝟏\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $0.7146$ $0.6315$ $\hat{K}=3$
Ray $n=1000,d=5,K=4$ ARI NMI CE $\boldsymbol{0.7577}$ $\boldsymbol{0.7541}$ $\boldsymbol{9.65\%}$ $0.7090$ $0.7439$ $11.90\%$ $0.4189$ $0.4847$ $31.40\%$ $0.5364$ $0.6175$ $37.70\%$ $0.6117$ $0.6595$ $21.10\%$ $0.6794$ $0.7013$ $13.10\%$ $0.2814$ $0.4245$ $\hat{K}=14$
Starbeam $n=780,d=2,K=2$ ARI NMI CE $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $0.4822$ $0.5072$ $15.26\%$ $0.4680$ $0.4982$ $15.77\%$ $0.5957$ $0.4671$ $11.28\%$ 0.0248-0.0248 $0.0379$ $39.23\%$ $0.9948$ $0.9859$ $0.13\%$ $0.2823$ $0.4351$ $\hat{K}=4$
Table II summarizes the clustering performance on the $12$ synthetic benchmarks. Overall, DLCC delivers consistently strong results and remains competitive across a broad range of challenges. In particular, DLCC attains perfect recovery on $7$ datasets, next to ACLR (an advanced graph-based method), and achieves favorable ARI/NMI with low CE on most of the remaining examples, illustrating its ability to handle both convex and non-convex structures as well as varying densities and close proximity.
Nevertheless, DLCC faces difficulties on Atom and on the spiral structures in Blend. As discussed in Section III-A, the local neighborhoods used to define local centers are expected to be approximately convex. This assumption may fail for manifold-like structures, where the neighborhood geometry around most points is non-convex. A representative failure case is Atom, in which the depth-defined local centers fail to adequately capture this geometry. A similar issue arises for the spiral components in Blend. This limitation of depth-defined local centers explains the observed performance degradation and highlights an inherent weakness of the current DLCC model.
V-D Real-world Datasets
As for the real-world benchmarks, we evaluate DLCC on $13$ datasets. Specifically, Iris, Seed, Wine, Parkinsons (Pa), Breast Cancer (BC), Segmentation (Seg), Optidigits, and Pendigits are obtained from the UCI Machine Learning Repository [64]. A highly imbalanced dataset, Anuran Calls (Anuran) is also included, whose cluster sizes ranging from as few as $68$ to as many as $3478$ [65]. We further include high-dimensional datasets ($d\geq 100$) with $n>d$, namely Yale B [44] and HAR [66], representing face images and human-activity signals, respectively. Among these, Yale B, Optidigits, and Pendigits are accessed via the R package PPCI [67]. In addition, we consider two microarray datasets, Leukemia [68] and SuCancer [69], which are characterized by extremely high dimensionality and small sample sizes. Overall, this benchmark suite is representative, spanning tabular, image, and time-series datasets, as well as a wide range of dimensionalities (low-dimensional to microarray-scale) and sample sizes.
For these real-world datasets, we adopt the max strategy for Pa, Anuran, Leukemia, and SuCancer due to severe class imbalance, and use the min strategy for all remaining datasets. Table III shows that DLCC is consistently competitive across the $13$ real-world benchmarks. It achieves the best performance on $6$ datasets and ranks within the top three on all datasets. DLCC also proves highly adaptable for high-dimensional data applications, attaining the strongest results on Yale B and HAR; on SuCancer it is still among the best, although the absolute performance is modest for all methods.
The table further indicates that competing methods are often dataset-dependent: some methods can be outstanding on particular datasets (e.g., EKM on Wine, R-MDPC on Optidigits, and ACLR on several benchmarks), yet no single method dominates uniformly. This is largely because different algorithms have different notions of what constitutes a “cluster” (e.g., centroid-based compactness, graph connectivity, or density/shape regularity), which may align well with some datasets but not with others. This variability underscores the practical value of DLCC as a robust alternative with strong performance across different real-world scenarios.
@@ V-D Real-world Datasets
TABLE III: Comparative analysis of selected clustering algorithms on real datasets.
Datasets Metric DLCC LW-k-means EKM R-MDPC GBCT ACLR DMS
Iris $n=150,d=4,K=3$ ARI NMI CE $\boldsymbol{0.8857}$ $\boldsymbol{0.8642}$ $\boldsymbol{4.00\%}$ $\boldsymbol{0.8857}$ $\boldsymbol{0.8642}$ $\boldsymbol{4.00\%}$ $0.6101$ $0.6526$ $17.33\%$ $\boldsymbol{0.8857}$ $\boldsymbol{0.8642}$ $\boldsymbol{4.00\%}$ $0.5638$ $0.7038$ $26.00\%$ $0.7592$ $0.7960$ $9.33\%$ $0.6956$ $0.7711$ $12.67\%$
Wine $n=178,d=13,K=3$ ARI NMI CE $0.8975$ $0.8759$ $3.37\%$ $0.8516$ $0.8267$ $5.06\%$ $\boldsymbol{0.9134}$ $\boldsymbol{0.8920}$ $\boldsymbol{2.81\%}$ $0.7269$ $0.7435$ $9.55\%$ $0.7033$ $0.7397$ $10.67\%$ $0.4041$ $0.4127$ $26.97\%$ $0.2748$ $0.3933$ $36.52\%$
Seed $n=210,d=7,K=3$ ARI NMI CE $0.7868$ $0.7409$ $\boldsymbol{7.62}\%$ $0.6687$ $0.6728$ $12.86\%$ $0.7715$ $0.7315$ $8.10\%$ $0.7289$ $0.6946$ $10.00\%$ $0.1614$ $0.3061$ $49.52\%$ $0.7132$ $0.6841$ $10.48\%$ $\boldsymbol{0.7870}$ $\boldsymbol{0.7423}$ $\boldsymbol{7.62}\%$
Pa $n=195,d=22,K=2$ ARI NMI CE $\boldsymbol{0.4226}$ $\boldsymbol{0.3265}$ $\boldsymbol{14.36}\%$ 0.0937-0.0937 $0.1189$ $44.10\%$ 0.0710-0.0710 $0.0558$ $31.79\%$ $0.1256$ $0.2776$ $31.79\%$ 0.0515-0.0515 $0.0360$ $29.23\%$ $0.2890$ $0.1965$ $17.44\%$ $0.0499$ $0.1767$ $\hat{K}=3$
BC $n=569,d=30,K=2$ ARI NMI CE $0.7361$ $0.6352$ $7.03\%$ $0.7177$ $0.6215$ $7.56\%$ $0.6466$ $0.5532$ $9.67\%$ $0.5175$ $0.5009$ $13.71\%$ 0.0302-0.0302 $0.0553$ $42.18\%$ $\boldsymbol{0.8241}$ $\boldsymbol{0.7324}$ $\boldsymbol{4.57\%}$ $0.0024$ $0.0188$ $37.08\%$
Seg $n=2086,d=18,K=7$ ARI NMI CE $\boldsymbol{0.6278}$ $\boldsymbol{0.6971}$ $\boldsymbol{21.86}\%$ $0.4782$ $0.6141$ $41.04\%$ $0.4862$ $0.6839$ $43.29\%$ $0.5832$ $0.7265$ $34.42\%$ $0.2407$ $0.5731$ $53.40\%$ $0.3829$ $0.5133$ $48.95\%$ $0.4981$ $0.6178$ $\hat{K}=8$
Yale B $n=2000,d=600,K=10$ ARI NMI CE $\boldsymbol{0.9846}$ $\boldsymbol{0.9851}$ $\boldsymbol{0.70\%}$ $0.8461$ $0.9306$ $13.25\%$ $0.1549$ $0.4090$ $60.20\%$ $0.7427$ $0.8765$ $17.35\%$ $0.1725$ $0.5722$ $61.10\%$ $0.8421$ $0.9097$ $14.65\%$ $0.5882$ $0.7744$ $33.80\%$
Optidigits $n=5620,d=64,K=10$ ARI NMI CE $0.8233$ $0.8394$ $8.83\%$ $0.4923$ $0.6486$ $35.96\%$ $0.0144$ $0.1290$ $82.99\%$ $\boldsymbol{0.8543}$ $\boldsymbol{0.8978}$ $\boldsymbol{8.75\%}$ $0.0003$ $0.0426$ $88.19\%$ $0.7326$ $0.7900$ $22.14\%$ $0.5671$ $0.6549$ $\hat{K}=11$
Pendigits $n=10992,d=16,K=10$ ARI NMI CE $0.6649$ $0.7433$ $18.31\%$ $0.5749$ $0.6912$ $27.85\%$ $0.5174$ $0.6484$ $30.04\%$ $0.6739$ $0.7840$ $19.17\%$ $0.0013$ $0.0754$ $86.14\%$ $\boldsymbol{0.7695}$ $\boldsymbol{0.8368}$ $\boldsymbol{12.50\%}$ $0.6696$ $0.7497$ $\hat{K}=17$
HAR $n=10299,d=561,K=6$ ARI NMI CE $\boldsymbol{0.6693}$ $\boldsymbol{0.7248}$ $\boldsymbol{17.74}\%$ $0.4494$ $0.5873$ $42.81\%$ $0.2932$ $0.4633$ $55.20\%$ $0.4995$ $0.6468$ $35.51\%$ $0.3279$ $0.5455$ $64.21\%$ $0.5994$ $0.6355$ $35.74\%$ $0.4075$ $0.5633$ $\hat{K}=9$
Anuran $n=7195,d=22,K=10$ ARI NMI CE $0.8592$ $0.7146$ $\hat{K}=11$ $0.5859$ $0.6551$ $34.37\%$ $0.5145$ $0.6229$ $58.21\%$ $0.7895$ $0.6489$ $25.70\%$ $0.0289$ $0.0727$ $50.02\%$ $\boldsymbol{0.8737}$ $\boldsymbol{0.7559}$ $\boldsymbol{24.04\%}$ $0.8296$ $0.6857$ $\hat{K}=9$
Leukemia $n=72,d=3571,K=2$ ARI NMI CE $0.8897$ $0.8056$ $2.78\%$ $0.8897$ $0.8056$ $2.78\%$ $0.2127$ $0.2178$ $26.39\%$ $0.7851$ $0.6790$ $5.56\%$ $0.0095$ $0.0036$ $34.72\%$ $\boldsymbol{0.9440}$ $\boldsymbol{0.8954}$ $\boldsymbol{1.39\%}$ $0.5905$ $0.5536$ $11.11\%$
SuCancer $n=174,d=7909,K=2$ ARI NMI CE $\boldsymbol{0.1154}$ $\boldsymbol{0.2317}$ $\boldsymbol{32.76}\%$ 0.0030-0.0030 $0.0041$ $47.70\%$ 0.0045-0.0045 $0.0006$ $48.28\%$ $\boldsymbol{0.1154}$ $\boldsymbol{0.2317}$ $\boldsymbol{32.76}\%$ 0.0022-0.0022 $0.0290$ $49.43\%$ $\boldsymbol{0.1154}$ $0.1808$ $\boldsymbol{32.76\%}$ $0.0169$ $0.0214$ $42.53\%$
VI Conclusions and Discussions